OpenStudy (anonymous):

What value of makes the equation true? 30/48 = 6/a A. 9 B. 30 C. 6 D. 225

TheSmartOne (thesmartone):

\[\frac{ 30 }{ 48 }=\frac{ 6 }{ a}\]

TheSmartOne (thesmartone):

Just cross multiply.

TheSmartOne (thesmartone):

\[30 a=6 \times 48\]

TheSmartOne (thesmartone):

\[a=\frac{ 6 \times 48 }{ 30 }\]

TheSmartOne (thesmartone):

So what does a=?

OpenStudy (confluxepic):

\[\huge \color{blue}{a=9.6}\]

TheSmartOne (thesmartone):

Correct ^^

OpenStudy (confluxepic):

@TheSmartOne showed all the correct work.

OpenStudy (anonymous):

I when I was working it out it came to that but I thought it was in-correct but ty for showing me how to do these problems ;)